# Splitting links by integer homology spheres

### v2

@@ 1. Introduction
A link in $S^{n+1}$ is the image of an embedding of a disjoint union of $(n-1)$-spheres, which we consider up to ambient isotopy. We recall that a link is called (smoothly or topologically) split if there is a (smooth or locally-flat) embedding of $S^{n}$ in $S^{n+1}$ that separates the link, i.e. each component of $S^{n+1}\setminus S^{n}$ contains at least one component of the link. More generally, given a closed $n$-manifold $Y$, we say that a link in $S^{n+1}$ is (smoothly or topologically) split by $Y$ if there exists a (smooth or locally-flat) embedding of $Y$ in $S^{n+1}$ that separates the link.
@@
We note that in the case $n=2$ (i.e. the case of classical links in $S^{3}$) the generalised notion of “split by $Y$” is not very useful, because every link with at least $2$ components is separated by a torus, which is the simplest possible surface after a 2-sphere. On the other hand, such a notion is highly non-trivial whenever $n\geq 3$. A simple homological argument provides an obstruction to being split by a rational homology sphere: if $L$ is split by a rational homology sphere, then the (single-variable) Alexander polynomial vanishes (see Remark 2.5). Homological arguments can also be used to produce examples of triples $(L,Y,Z)$, where $Y$ and $Z$ are rational homology spheres and $L$ is a link split by $Y$ but not by $Z$ (see Remark 2.4; these examples have $H_{1}(Y)\neq H_{1}(Z)$). Exhibiting such triples where $Y$ and $Z$ are integer homology spheres, as in Theorem 1.1, is more subtle, and requires arguments that go beyond homology.
@@ Acknowledgements
We would like to thank Peter Kronheimer for raising the question and Maggie Miller for the talk on [HKM23] which inspired it. We are also indebted to Michel Boileau for his suggestions and for pointing us to Rong’s paper [Ron95]. Finally, we thank Clayton McDonald for his comments on a draft of this paper and Danny Ruberman for enlightening conversations about covers. MG thanks the Rényi Institute for their hospitality. MG was partially supported by the Étoile Montante project PSyCo of the Region Pays de la Loire. MM was partially supported by NKFIH grant OTKA K146401.
@@ 2. The construction
We define a link $L_{K}\subset S^{n+1}$ by taking the disjoint union of two push-offs of the binding $S$, one in each normal direction to $Y_{K}$. Call these two $(n-1)$-spheres $S^{+}_{K}\subset W^{+}_{K}$ and $S^{-}_{K}\subset W^{-}_{K}$. Note that in this way, by construction, the link is split by $Y_{K}$.
@@
×$\mathbb{R}\times Y\to\Sigma_{2}(S^{n+1},L)$
obtained by unwinding the mapping torus is in fact the maximal Abelian cover of $\Sigma_{2}(S^{n+1},L)$. Again, since $H_{1}(Z;\mathbb{Z})=0$, we can lift $Z$ to ×$\mathbb{R}\times Y$, again as a homological section. By choosing two parallel lifts of $Y$ on either sides of $Z$ we obtain an identity cobordism $I\times Y$ which is split by $Z$. One of the two components of $(I\times Y)\setminus Z$ gives a cobordism from $Y$ to $Z$, and the other part is the left inverse cobordism. ∎
@@
We can then follow the argument of Lemma 2.3. First one shows that $Z$ lifts from $S^{n}\setminus L$ to $\Sigma_{2}(L)\setminus\pi^{-1}(L)$, using the fact that the two-fold cover is Abelian, and that the torsion $\mathbb{Z}$-module $H_{1}(Z)$ must map trivially to the free $\mathbb{Z}$-module $H_{1}(S^{n}\setminus L)$. Then, the fact that $\Sigma_{2}(L)$ is a mapping torus over $Y$ implies via a Mayer–Vietoris sequence that $H_{1}(\Sigma_{2}(L))\cong\mathbb{Z}\oplus Q$, where $Q$ is a quotient of $H_{1}(Y)$. Thus, the map $H_{1}(Z)\to H_{1}(\Sigma_{2}(L))$ vanishes, and thus $Z$ lifts to the Abelian cover $I\times Y$. Moreover, $Z$ is homologically primitive in $I\times Y$, by the same argument as in Lemma 2.3, so $[Y]=\pm[Z]$ in $H_{3}(I\times Y)\cong\mathbb{Z}$. Thus, restricting the projection gives a degree-one map from $Z$ to $Y$, which must induce a surjection in homology, and this gives the desired contradiction.
@@
If one wants to construct links that are not split by any rational homology sphere, then there is a simple obstruction coming from the single-variable Alexander polynomial: if $L$ is split by $Y$ and $b_{1}(Y)=0$, then $\Delta_{L}(t)=0$. To see this, note that $Y$ lifts to the infinite cyclic cover $\tilde{X}$ of $L$ associated to the morphism that sends each meridian $\mu_{i}$ of $L$ to a generator of $\mathbb{Z}$ (here we are taking $L$ with an orientation), and so the preimage of $Y$ to $\tilde{X}$ is $\coprod_{m\in\mathbb{Z}}t^{m}\,Y_{0}$. Consider a loop $\gamma$ in $S^{4}$ in the homology class $[\mu_{1}]-[\mu_{2}]$ that intersects $Y$ geometrically twice. See Figure 2. $\gamma$ also lifts to a loop $\gamma_{0}$ in $\tilde{X}$, and the pairing satisfies:
@@
Thus, the $\mathbb{Z}[t^{\pm 1}]$-module $H_{1}(\tilde{X};\mathbb{Z})$ is not torsion, and therefore its first Alexander ideal vanishes.
@@ 3. Higher-dimensional links
By Lemma 2.3 there is an invertible topological cobordism from $Y$ to $S^{n}$, or, equivalently, there is a locally-flat embedding of $S^{n}$ in ×$\mathbb{R}\times Y$ as a homological section. Composing with the projection, this would give a degree-1 map from $S^{n}$ to $Y$. A degree-1 map always induces an epimorphism on $\pi_{1}$, but this is impossible under the assumption that $\pi_{1}(Y)\neq 1$. ∎
@@
Due to the presence of the hypothesis $\pi_{1}(Y)\neq 1$ in the statement of Proposition 3.1, the reader may wonder if it is possible to find a 2-component link in $S^{n+1}$ which is split by an $n$-dimensional homotopy sphere, but not by $S^{n}$.
When $n\neq 4$ the answer is negative, because it is impossible to even embed an $n$-dimensional exotic sphere $Y$ in $S^{n+1}$. Indeed, an embedding $Y\hookrightarrow S^{n+1}$ would give an $h$-cobordism between $Y$ and $S^{n}$ (after removing a small ball from $S^{n+1}$), hence showing that $Y$ is diffeomorphic to $S^{n}$ [Sma62]. (Note that $n\geq 7$ because $Y$ is assumed to be exotic.)
When $n=4$ it is unknown whether an exotic $S^{4}$ exists, and the $h$-cobordism theorem does not hold, so in principle it is possible to split a link with an exotic $S^{4}$. Doing so would require refuting the smooth 4-dimensional Poincaré conjecture, so we will save this problem for another day.
@@
We define a knot $K$ in $S^{n}$ by taking the 00-spun knot $n-3$ times:
@@
From the 00-spinning construction, it is straightforward to check that if $J$ is a knot whose double branched cover is an integer homology sphere, then so is $Sp_{0}(J)$. Thus, iterating this reasoning, we have that $\Sigma_{2}(S^{n},K)$ is an integer homology sphere. Moreover, since 00-spinning preserves the knot group, we have that $\pi_{1}(S^{n}\setminus K)\cong\pi_{1}(S^{3}\setminus J)$, and therefore $\pi_{1}(\Sigma_{2}(S^{n},K))\cong\pi_{1}(\Sigma_{2}(S^{3},J))$, being their unique index-$2$ subgroups: thus, $\Sigma_{2}(S^{n},K)$ is also an integer homology sphere with $\pi_{1}(\Sigma_{2}(S^{n},K))\neq 1$.
@@
4. Locally-flatly splitting in dimension $4$
@@
By Lemma 2.3 there is an invertible cobordism from $Z\#\overline{Z}$ to $Y\#\overline{Y}$, and therefore a topological embedding of $Y\#\overline{Y}$ in ×$\mathbb{R}\times(Z\#\overline{Z})$ such that post-composing with the projection ×$\mathbb{R}\times(Z\#\overline{Z})\to Z\#\overline{Z}$ yields a continuous degree-$1$ map
@@
Since $Z$ is a Seifert fibred space with infinite fundamental group, $\pi_{3}(Z)=0$, and therefore $g_{A}$ is null-homotopic, and in particular it has degree 00. It follows that also $f_{A}$ has degree 00 (since $A$ is the preimage of $Z\setminus B^{3}\subset Z$, and we can compute the degree by counting the points in the preimage of a generic point).
Since the map $f$ has degree $1$, and its restriction to all these copies of $S^{2}\times[0,1]$ has degree 00, the preimage of $Z\setminus B^{3}$ cannot be contained in a union of these copies, and in particular it must contain, say, $Y\setminus B^{3}\subset Y\#\overline{Y}$. It follows that the restriction to $Y\setminus B^{3}$ has degree $1$ onto $Z\setminus B^{3}$. We can now extend the restriction of $f$ to $Y\setminus B^{3}$ to a degree-$1$ map $g\colon Y\to Z$, but no such map exists by work of Rong [Ron95, Theorem 3.2], as explained above.
@@
5. Smoothly splitting in dimension $4$
@@
The following definition is a compact way of referring to integer homology spheres with $\operatorname{HF}_{\rm red}$ of a special form, which is convenient for a set of examples that we will study. We say that a homology 3-sphere $Y$ is of size $\ell\geq 2$ if:
@@
  • there are exactly two towers of length $1$ in $\operatorname{HF}^{-}(\Sigma_{\ell})$, supported in Maslov gradings $-\ell(\ell+1)$ and $-(\ell-1)\ell$;
@@
  • the closest grading supporting the generator of a finite tower in $\operatorname{HF}^{-}(\Sigma_{\ell})$ to $-\ell(\ell+1)$ or $-(\ell-1)\ell$ is $-(\ell-2)(\ell-1)$.
@@
(Note that there is a degree shift by $+1$ in $\operatorname{HF}_{\textrm{red}}$ when changing the orientation, and a degree shift by $-1$ when applying the Tor functor in the tensor product, and they cancel out.)
@@ Appendix A Heegaard Floer homology
If $Y$ is an integer homology sphere, one of the groups associated to $Y$ is called $\operatorname{HF}^{-}(Y)$: this is a finitely generated $\mathbb{Z}$-graded $\mathbb{F}[U]$-module, where $U$ is a formal variable of degree $-2$. The degree is called the Maslov grading. It sits into an exact sequence of $U$-equivariant degree-00 maps [^2] A word of caution here. There is also an $\operatorname{HF}^{+}_{\mathrm{red}}$, which is a quotient of $\operatorname{HF}^{+}$, and it differs from $\operatorname{HF}^{-}$ by a grading shift: $\operatorname{HF}^{-}_{\rm red}\cong\operatorname{HF}^{+}_{\rm red}[+1]$ as graded $\mathbb{F}[U]$-modules. When we use $\operatorname{HF}_{\rm red}$ we refer to either of them, but we assume that the choice we make is coherent.:
@@
We denote with $\operatorname{HF}^{-}_{\mathrm{even}}(Y)$ (resp. $\operatorname{HF}^{-}_{\mathrm{red,even}}(Y)$) the even-degree summand of $\operatorname{HF}^{-}(Y)$ (resp. $\operatorname{HF}^{-}_{\mathrm{red}}(Y)$). Since $U$ has degree $-2$, $\operatorname{HF}^{-}_{\mathrm{even}}(Y)$ is an $\mathbb{F}[U]$-submodule of $\operatorname{HF}^{-}(Y)$.
@@
which has degree 00, which descends to an analogous homomorphism in $\operatorname{HF}^{-}_{\mathrm{red}}$. This association is functorial: if $W^{\prime}$ is another rational homology cobordism from $Y^{\prime}$ to $Y^{\prime\prime}$, then the composition of the associated maps is the map associated to the cobordism $W^{\prime\prime}$ obtained by stacking $W$ and $W^{\prime}$. In other words,

### v3

@@ 1. Introduction
A link in $S^{n+1}$ is the image of an embedding of a disjoint union of (n1)(n-1)-spheres, which we consider up to ambient isotopy. We recall that a link is called (smoothly or topologically) split if there is a (smooth or locally-flat) embedding of $S^{n}$ in $S^{n+1}$ that separates the link, i.e. each component of $S^{n+1}\setminus S^{n}$ contains at least one component of the link. More generally, given a closed $n$-manifold $Y$, we say that a link in $S^{n+1}$ is (smoothly or topologically) split by $Y$ if there exists a (smooth or locally-flat) embedding of $Y$ in $S^{n+1}$ that separates the link.
@@
We note that in the case $n=2$ (i.e. the case of classical links in $S^{3}$) the generalised notion of “split by YY” is not very useful, because every link with at least $2$ components is separated by a torus, which is the simplest possible surface after a 2-sphere. On the other hand, such a notion is highly non-trivial whenever $n\geq 3$. A simple homological argument provides an obstruction to being split by a rational homology sphere: if $L$ is split by a rational homology sphere, then the (single-variable) Alexander polynomial vanishes (see Remark 2.5). Homological arguments can also be used to produce examples of triples $(L,Y,Z)$, where $Y$ and $Z$ are rational homology spheres and $L$ is a link split by $Y$ but not by $Z$ (see Remark 2.4; these examples have $H_{1}(Y)\neq H_{1}(Z)$). Exhibiting such triples where $Y$ and $Z$ are integer homology spheres, as in Theorem 1.1, is more subtle, and requires arguments that go beyond homology.
@@ Acknowledgements
We would like to thank Peter Kronheimer for raising the question and Maggie Miller for the talk on [HKM23] which inspired it. We are also indebted to Michel Boileau for his suggestions and for pointing us to Rong’s paper [Ron95]. We thank Clayton McDonald for his comments on a draft of this paper and Danny Ruberman for enlightening conversations about covers. Finally, we thank the anonymous referee for their insightful comments. MG thanks the Rényi Institute for their hospitality. MG was partially supported by the Étoile Montante project PSyCo of the Region Pays de la Loire. MM was partially supported by NKFIH grant OTKA K146401.
@@ 2. The construction
Here and below, $m(K)$ denotes the mirror of the knot $K$, and therefore $\Sigma_{2}(S^{n},m(K))$ is diffeomorphic to $\Sigma_{2}(S^{n},K)$.
@@
We define a link $L_{K}\subset S^{n+1}$ by taking the disjoint union of two push-offs of the binding $S$, one in each normal direction to $Y_{K}$. Call these two (n1)(n-1)-spheres $S^{+}_{K}\subset W^{+}_{K}$ and $S^{-}_{K}\subset W^{-}_{K}$. Note that in this way, by construction, the link is split by $Y_{K}$.
@@
$\mathbb{R}\times Y\to\Sigma_{2}(S^{n+1},L)$
obtained by unwinding the mapping torus is in fact the maximal Abelian cover of $\Sigma_{2}(S^{n+1},L)$. Again, since $H_{1}(Z;\mathbb{Z})=0$, we can lift $Z$ to $\mathbb{R}\times Y$, again as a homological section. By choosing two parallel lifts of $Y$ on either sides of $Z$ we obtain an identity cobordism $I\times Y$ which is split by $Z$. One of the two components of $(I\times Y)\setminus Z$ gives a cobordism from $Y$ to $Z$, and the other part is the left inverse cobordism. ∎
@@
We can then follow the argument of Lemma 2.3. First one shows that $Z$ lifts from $S^{n}\setminus L$ to $\Sigma_{2}(L)\setminus\pi^{-1}(L)$, using the fact that the two-fold cover is Abelian, and that the torsion $\mathbb{Z}$-module $H_{1}(Z)$ must map trivially to the free $\mathbb{Z}$-module $H_{1}(S^{n}\setminus L)$. Then, the fact that $\Sigma_{2}(L)$ is a mapping torus over $Y$ implies via a Mayer–Vietoris sequence that $H_{1}(\Sigma_{2}(L))\cong\mathbb{Z}\oplus Q$, where $Q$ is a quotient of $H_{1}(Y)$. Thus, the map $H_{1}(Z)\to H_{1}(\Sigma_{2}(L))$ vanishes, and thus $Z$ lifts to the Abelian cover $I\times Y$. Moreover, $Z$ is homologically primitive in $I\times Y$, by the same argument as in Lemma 2.3, so $Y]=\pm[Z]$ in $H_{3}(I\times Y)\cong\mathbb{Z}$. Thus, restricting the projection gives a degree-one map from $Z$ to $Y$, which must induce a surjection in homology, and this gives the desired contradiction.
@@
If one wants to construct links that are not split by any rational homology sphere, then there is a simple obstruction coming from the single-variable Alexander polynomial: if $L$ is split by $Y$ and $b_{1}(Y)=0$, then $\Delta_{L}(t)=0$. To see this, note that $Y$ lifts to the infinite cyclic cover $\tilde{X}$ of $L$ associated to the morphism that sends each meridian $\mu_{i}$ of $L$ to a generator of $\mathbb{Z}$ (here we are taking $L$ with an orientation), and so the preimage of $Y$ to $\tilde{X}$ is $\coprod_{m\in\mathbb{Z}}t^{m}\,Y_{0}$. Consider a loop $\gamma$ in $S^{4}$ in the homology class $\mu_{1}]-[\mu_{2}]$ that intersects $Y$ geometrically twice. See Figure 2. $\gamma$ also lifts to a loop $\gamma_{0}$ in $\tilde{X}$, and the pairing satisfies:
@@
[…]
@@ 3. Higher-dimensional links
By Lemma 2.3 there is an invertible topological cobordism from $Y$ to $S^{n}$, or, equivalently, there is a locally-flat embedding of $S^{n}$ in $\mathbb{R}\times Y$ as a homological section. Composing with the projection, this would give a degree-1 map from $S^{n}$ to $Y$. A degree-1 map always induces an epimorphism on $\pi_{1}$, but this is impossible under the assumption that $\pi_{1}(Y)\neq 1$. ∎
@@
Due to the presence of the hypothesis $\pi_{1}(Y)\neq 1$ in the statement of Proposition 3.1, the reader may wonder if it is possible to find a 2-component link in $S^{n+1}$ which is split by an $n$-dimensional homotopy sphere, but not by $S^{n}$. The answer is always negative.
When $n\neq 4$ it is impossible to even embed an $n$-dimensional exotic sphere $Y$ in $S^{n+1}$. Indeed, an embedding $Y\hookrightarrow S^{n+1}$ would give an $h$-cobordism between $Y$ and $S^{n}$ (after removing a small ball from $S^{n+1}$), hence showing that $Y$ is diffeomorphic to $S^{n}$ [Sma62]. (Note that $n\geq 7$ because $Y$ is assumed to be exotic.)
When $n=4$, even if the existence of exotic 4-spheres is open, one can still show that a homotopy sphere $Y^{4}$ splits a link $L$ if and only if the standard sphere $S^{4}$ does. This can be deduced from the fact that there is a unique smooth $h$-cobordism between any two homotopy 4-spheres $Y$ and $Y^{\prime}$, up to diffeomorphism rel boundary [Law88, Kre01]. Given the unique $h$-cobordism $W\colon Y\to Y^{\prime}$ and its inverse $W\colon Y^{\prime}\to Y$, the composition $(-W)\circ W$ must be diffeomorphic to $I\times Y$, by uniqueness. It follows that $Y^{\prime}$ sits in $I\times Y$ as a homological section. In other words, a regular neighbourhood $\nu(Y)$ of a homotopy 4-sphere $Y\subset S^{5}$ contains every other homotopy 4-sphere $Y^{\prime}$ as a homological section, so $L$ is split by a homotopy 4-sphere if and only if it is split by all of them.
@@
We define a knot $K$ in $S^{n}$ by taking the 0-spun knot n3n-3 times:
@@
From the 0-spinning construction, it is straightforward to check that if $J$ is a knot whose double branched cover is an integer homology sphere, then so is $Sp_{0}(J)$. Thus, iterating this reasoning, we have that $\Sigma_{2}(S^{n},K)$ is an integer homology sphere. Moreover, since 0-spinning preserves the knot group, we have that $\pi_{1}(S^{n}\setminus K)\cong\pi_{1}(S^{3}\setminus J)$, and therefore $\pi_{1}(\Sigma_{2}(S^{n},K))\cong\pi_{1}(\Sigma_{2}(S^{3},J))$, being their unique index-$2$ subgroups: thus, $\Sigma_{2}(S^{n},K)$ is also an integer homology sphere with $\pi_{1}(\Sigma_{2}(S^{n},K))\neq 1$.
@@
4. Locally-flatly splitting in dimension 44
@@
By Lemma 2.3 there is an invertible cobordism from $Z\#\overline{Z}$ to $Y\#\overline{Y}$, and therefore a topological embedding of $Y\#\overline{Y}$ in $\mathbb{R}\times(Z\#\overline{Z})$ such that post-composing with the projection $\mathbb{R}\times(Z\#\overline{Z})\to Z\#\overline{Z}$ yields a continuous degree-$1$ map
@@
Since $Z$ is a Seifert fibred space with infinite fundamental group, $\pi_{3}(Z)=0$, and therefore $g_{A}$ is null-homotopic, and in particular it has degree 0. It follows that also $f_{A}$ has degree 0 (since $A$ is the preimage of $Z\setminus B^{3}\subset Z$, and we can compute the degree by counting the points in the preimage of a generic point).
Since the map $f$ has degree $1$, and its restriction to all these copies of $S^{2}\times[0,1]$ has degree 0, the preimage of $Z\setminus B^{3}$ cannot be contained in a union of these copies, and in particular it must contain, say, $Y\setminus B^{3}\subset Y\#\overline{Y}$. It follows that the restriction to $Y\setminus B^{3}$ has degree $1$ onto $Z\setminus B^{3}$. We can now extend the restriction of $f$ to $Y\setminus B^{3}$ to a degree-$1$ map $g\colon Y\to Z$, but no such map exists by work of Rong [Ron95, Theorem 3.2], as explained above.
@@
5. Smoothly splitting in dimension 44
@@
The following definition is a compact way of referring to integer homology spheres with $\operatorname{HF}_{\rm red}$ of a special form, which is convenient for the set of examples that we will study here. We say that a homology 3-sphere $Y$ is of size $\ell\geq 2$ if:
@@
  • there are exactly two towers of length $1$ in $\operatorname{HF}^{-}(\Sigma_{\ell})$, supported in Maslov gradings $\ell(\ell+1)$ and $(\ell-1)\ell$;
@@
  • the closest grading supporting the generator of a finite tower in $\operatorname{HF}^{-}(\Sigma_{\ell})$ to $\ell(\ell+1)$ or $(\ell-1)\ell$ is $(\ell-2)(\ell-1)$.
@@
(Note that there is a degree shift by +1+1 in $\operatorname{HF}_{\textrm{red}}$ when changing the orientation, and a degree shift by 1-1 when applying the Tor functor in the tensor product, and they cancel out.)
@@ Appendix A Heegaard Floer homology
If $Y$ is an integer homology sphere, one of the groups associated to $Y$ is called $\operatorname{HF}^{-}(Y)$: this is a finitely generated $\mathbb{Z}$-graded $\mathbb{F}[U]$-module, where $U$ is a formal variable of degree 2-2. The degree is called the Maslov grading. It sits into an exact sequence of $U$-equivariant degree-0 maps [^2] A word of caution here. There is also an $\operatorname{HF}^{+}_{\mathrm{red}}$, which is a quotient of $\operatorname{HF}^{+}$, and it differs from $\operatorname{HF}^{-}$ by a grading shift: $\operatorname{HF}^{-}_{\rm red}\cong\operatorname{HF}^{+}_{\rm red}[+1]$ as graded $\mathbb{F}[U]$-modules. When we use $\operatorname{HF}_{\rm red}$ we refer to either of them, but we assume that the choice we make is coherent.:
@@
We denote with $\operatorname{HF}^{-}_{\mathrm{even}}(Y)$ (resp. $\operatorname{HF}^{-}_{\mathrm{red,even}}(Y)$) the even-degree summand of $\operatorname{HF}^{-}(Y)$ (resp. $\operatorname{HF}^{-}_{\mathrm{red}}(Y)$). Since $U$ has degree 2-2, $\operatorname{HF}^{-}_{\mathrm{even}}(Y)$ is an $\mathbb{F}[U]$-submodule of $\operatorname{HF}^{-}(Y)$.
@@
which has degree 0, which descends to an analogous homomorphism in $\operatorname{HF}^{-}_{\mathrm{red}}$. This association is functorial: if $W^{\prime}$ is another rational homology cobordism from $Y^{\prime}$ to $Y^{\prime\prime}$, then the composition of the associated maps is the map associated to the cobordism $W^{\prime\prime}$ obtained by stacking $W$ and $W^{\prime}$. In other words,